# Random coverage from within with variable radii, and Johnson-Mehl cover timesthanks: Supported by EPSRC grant EP/T028653/1

### v2

@@ Abstract
Given a compact planar region $A$, let $\tau_{A}$ be the (random) time it takes for $A$ to be fully covered by a spatial birth-growth process in $A$ with seeds arriving as a unit-intensity Poisson point process in $A\times[0,\infty)$, where upon arrival each seed grows at unit rate in all directions. We show that if $\partial A$ is smooth or polygonal then $\mathbb{P}[\pi\tau_{sA}^{3}-6\log s-4\log\log s\leq x]$ tends to $\exp(-(\frac{81}{4\pi})^{1/3}|A|e^{-x/3}\linebreak-(\frac{9}{2\pi^{2}})^{1/3}|\partial A|e^{-x/6})$ in the large-$s$ limit; the second term in the exponent is due to boundary effects, the importance of which was not recognized in earlier work on this model. We present similar results in higher dimensions (where boundary effects dominate). These results are derived using new results on the asymptotic probability of covering $A$ with a high-intensity spherical Poisson Boolean model restricted to $A$ with grains having iid small random radii, which generalize recent work of the first author that dealt only with grains of deterministic radius.
@@ 1 Introduction
This model dates back at least to work of Kolmogorov and of Johnson and Mehl in the 1930s on modelling crytallization processes; other applications include growth of surface film on metals, and more recent applications in neurobiology are described in [2]. For further discussion and references, see also [7, 8].
A natural variant is the restricted Johnson-Mehl tessellation, which we define as follows. Given a specified compact region $A\subset\mathbb{R}^{d}$, we allow only for seeds that arrive inside the region $A$, that is, the seeds are generated by a Poisson process in $A\times[0,\infty)$. Cells grow by the same rules as described before, and the restrictions of the cells to $A$ ultimately tessellate $A$.
@@
In [1, page 893], Chiu rightly distinguishes between the J-M and restricted J-M models. However, he goes on to assert that ‘all theorems in this paper are valid for both models’ because ‘almost surely edge effects do not play a role in the limiting behaviour of $T_{L}$’ (Chiu’s $T_{L}$ is equivalent to our $\tau_{L}$). In other words, Chiu seems to be asserting that $\tau_{L}$ and $\tilde{\tau}_{L}$ have the same limiting distribution.
It is our contention that this assertion is incorrect. In dimensions $d\geq 2$, the limiting behaviour of $T_{\rho}$ is different from that of $\tilde{T}_{\rho}$, and the limiting behaviour of $\tau_{L}$ is different from that of $\tilde{\tau}_{L}$; in other words, edge effects do play a role. We justify this assertion with results identifying the limiting distribution of $T_{\rho}$ (suitably scaled and centred) and showing it is different from that of $\tilde{T}_{\rho}$; our limiting result for $\tilde{T}_{\rho}$ is consistent with the result in [1]. Likewise we show that the limiting distribution of $\tau_{L}$ (suitably scaled and centred) is different from that of $\tilde{\tau}_{L}$. Our results for $T_{\rho}$ and $\tau_{L}$ apply when $A$ is polygonal in $d=2$, or when it has a smooth (more precisely, $C^{2}$) boundary for general $d\geq 2$.
When $A=[0,1]^{d}$, we do not provide detailed limiting distributions for $T_{\rho}$ and $\tau_{L}$ except in the case $d=2$. To give a detailed limit distribution when $d\geq 3$ would require careful consideration of all faces of all dimensions and is beyond the scope of this paper. However, the time to cover all the $(d-1)$-dimensional faces will be a lower bound for the actual cover time so boundary effects can certainly not be neglected in this case either.
@@ 2 Statement of results
[…]
@@
We say that $D$ has $C^{2}$ boundary (for short: $\partial D\in C^{2}$) if for each $x\in\partial D$ there exists a neighbourhood $U$ of $x$ and a real-valued function $f$ that is defined on an open set in $\mathbb{R}^{d-1}$ and twice continuously differentiable, such that $\partial D\cap U$, after a rotation, is the graph of the function $f$. We say that $\partial D\in C^{1,1}$ (a weaker condition) if for each $x$ the function $f$ satisfies only that $f$ is contiuously differentiable with Lipschitz first order partial derivatives.
2.1 Results for the spherical Poisson Boolean model
As discsussed in the introduction, we shall derive our results for the J-M model from results on coverage by the restricted SPBM, which are of independent interest and which we now state. Given a nonnegative random variable $Y$ and given $n>0$, suppose we have a collection of balls of independent random radii with the distribution of $Y$, centred on the points of a homogeneous Poisson point process in $A$ of intensity $n$.
Theorem 2.1.
$\mathbb{P}[A\subset Z_{A,k}(n,r_{n}Y)]\to\exp\Big{(}-\Big{(}\frac{(\mathbb{E}\,[Y])^{2}}{\mathbb{E}\,[Y^{2}]}\Big{)}{\bf 1}_{\{k=1\}}|A|e^{-\beta}-\Big{(}\frac{c_{2,k}\mathbb{E}\,[Y]|\partial A|}{(\mathbb{E}\,[Y^{2}])^{1/2}}\Big{)}e^{-\beta/2}\Big{)},$ (2.1)
where we set $c_{2,k}:=\pi^{-1/2}(1/2)^{k-1}/(k-1)!$.
@@ 2.1 Results for the spherical Poisson Boolean model
$c_{d}:=\frac{1}{d!}\left(\frac{\sqrt{\pi}\;\Gamma(1+d/2)}{\Gamma((d+1)/2)}\right)^{d-1}.$ (2.2)
Note that $c_{1}=c_{2}=1$, and $c_{3}=3\pi^{2}/32$. Moreover, using Stirling’s formula one can show that $c_{d}^{1/d}\sim e(\pi/(2d))^{1/2}$ as $d\to\infty$. The constant $c_{d}$ is denoted $\psi_{d}$ in [1, page 894], while $\omega_{d}$ is denoted $\theta_{d}$ in [10]. Later it will be useful to have $\omega_{0}$ defined: we set $\omega_{0}:=1$.
Our result for $A$ with $\partial A\in C^{2}$ involves a further constant $c_{d,k}$ defined (as in [10]) for each $d,k\in\mathbb{N}$ with $d\geq 2$ by
$c_{d,k}$ $:=c_{d-1}\omega_{d}^{2-d-1/d}\omega_{d-1}^{2d-3}\omega_{d-2}^{1-d}(1-1/d)^{d+k-3+1/d}2^{-1+1/d}/(k-1)!.$ (2.3)
It can be checked that
$c_{d,1}=((d-1)!)^{-1}2^{1-d}(d-1)^{d-2+1/d}\pi^{(d/2)-1}\Gamma\left(\frac{d+1}{2}\right)^{1-d}\Gamma\left(\frac{d}{2}\right)^{d-1+1/d},$ (2.4)
and $c_{d,k}=c_{d,1}(1-1/d)^{k-1}/(k-1)!$. By (2.3) we have $c_{2,1}=\pi^{-1/2}$ and $c_{3,1}=2^{-4}\pi^{5/3}$. In the right hand side of (2.4) the first factor of $((d-1)!)^{-1}$ was given, incorrectly, as $(d!)^{-1}$ in [10].
Theorem 2.2.
$n\omega_{d}r_{n}^{d}\mathbb{E}\,[Y^{d}]-(2-2/d)\log n-2(d+k-3+1/d)\log\log n\to\beta$ (2.5)
as $n\to\infty$. If $d=2$ then (2.1) holds, while if $d\geq 3$ then
$\lim_{n\to\infty}\mathbb{P}[A\subset Z_{A,k}(n,r_{n}Y)]=\exp\Big{(}-c_{d,k}\Big{(}\frac{(\mathbb{E}\,[Y^{d-1}])^{d-1}}{(\mathbb{E}\,[Y^{d}])^{d-2+1/d}}\Big{)}|\partial A|e^{-\beta/2}\Big{)}.$ (2.6)
Remark 2.3.
Taking $Y\equiv 1$ in these results gives [10, Theorem 3.2] and [10, Theorem 3.1]. [10] also includes parallel results (in the case $Y\equiv 1$) for a situation where the number of grains in the spherical Boolean model is a large deterministic constant rather than a Poisson variable. We expect that a result along these lines could be given for general $Y$ using the methods of this paper.
Remark 2.4.
The condition $\overline{A^{o}}=A$ should have been included in the statement of [10, Theorem 3.1]. For example, if $d=2$ and $A$ is the union of a disk (filled in) and a circle (not filled in, and disjoint from the disk) then that theorem does not apply.
Remark 2.5.
Remark 2.6.
If $\mathbb{E}\,[Y^{d}]=\infty$ then $\mathbb{P}[A\subset Z_{\mathbb{R}^{d},k}(n,rY)]=1$ for any $k\in\mathbb{N}$ and any $r>0$; see e.g. [6, Theorem 16.4] for the case $k=1$. However, one could still look for a sequence $(r_{n})$ such that $\mathbb{P}[A\subset Z_{A,1}(n,r_{n}Y)]$ converges to a nontrivial limit. It seems plausible in this case that $A$ is most likely to be covered (if at all) by a small number of balls. This looks like an interesting problem that lies beyond the scope of this paper.
Remark 2.7.
A series of videos illustrating the SPBM and Remark 2.6 can be viewed at this link [^2] https://www.youtube.com/playlist?list=PLiaV5rk6Gk7qEXpLOU7FSvN4b8dy1_GJn. There are three videos, with $n=1000$, $n=10000$ and $n=20000$ respectively. Each video has a Poisson point process of intensity $n$ inside $[0,1]^{2}$ with points $\{p_{1},\dots,p_{N_{n}}\}$. The frame at time $t$ is an image of a tessellation corresponding to a restricted SPBM, $Z_{A,1}(n,r_{t}Y_{t})$, where $A=[0,1]^{2}$, $Y_{t}$ is Pareto distributed with shape parameter $\alpha(t)$ and $r_{t}$ is chosen so that $A\subset Z_{A,1}(n,r_{t}Y_{t})$. There are $N_{n}$ cells in the frame: the pixel with centre $x\in[0,1]^{2}$ is assigned to a cell based on $\mathrm{argmin}_{i}\inf\{r\geq 0:x\in B(p_{i},rY_{t,i})\}$, where $(Y_{t,i})_{1\leq i\leq N_{n}}$ are independent copies of $Y_{t}$. In other words, for $t$ fixed the tessellation is generated by a growth process for which all seeds are born at time 0, growth rates are random and distributed like $Y_{t}$, and the growing cells are interpenetrable.
The videos after the point where $\mathbb{E}\,[Y_{t}^{2}]=\infty$ illustrate Remark 2.6. In the period when $\mathbb{E}\,[Y_{t}^{2}]=\infty$ but $\mathbb{E}\,[Y_{t}^{7/4}]<\infty$, there is not yet a single cell containing more than half the area of $[0,1]^{2}$, indicating that there may be a non-trivial period where several balls are required to cover $[0,1]^{2}$.
@@ 2.2 Results for the Johnson-Mehl model
$T_{\rho}:=\inf\left\{t>0:A\subset\cup_{(x,s)\in{\cal H}_{\rho}\cap(A\times[0,t])}B(x,t-s)\right\};$ (2.7)
$\tilde{T}_{\rho}:=\inf\left\{t>0:A\subset\cup_{(x,s)\in{\cal H}_{\rho}\cap(\mathbb{R}^{d}\times[0,t])}B(x,t-s)\right\}.$ (2.8)
@@
$\tau_{L}:=\inf\left\{t>0:A_{L}\subset\cup_{(x,s)\in{\cal H}_{1}\cap(A_{L}\times[0,t])}B(x,t-s)\right\};$ (2.9)
$\tilde{\tau}_{L}:=\inf\left\{t>0:A_{L}\subset\cup_{(x,s)\in{\cal H}_{1}\cap(\mathbb{R}^{d}\times[0,t])}B(x,t-s)\right\}.$ (2.10)
@@
Our results in this subsection are concerned with the asymptotic distribution of the cover times $T_{\rho}$ and $\tilde{T}_{\rho}$ as $\rho\to\infty$, and the asymptotic distribution of $\tau_{L}$ and $\tilde{\tau}_{L}$ as $L\to\infty$. We start with a result in general $d$ for the unrestricted cover time $\tilde{T}_{\rho}$ (and for $\tilde{\tau}_{L})$. This is simpler to deal with than $T_{\rho}$ because boundary effects are avoided.
@@
Suppose $A\subset\mathbb{R}^{d}$ is compact and Riemann measurable with $|A|>0$. Let $\beta\in\mathbb{R}$. Then as $\rho\to\infty$,
$\mathbb{P}\left[\omega_{d}\rho\tilde{T}_{\rho}^{d+1}-d\log\rho-d^{2}\log\log\rho\leq\beta\right]\to\exp(-c_{d}(d^{d}\omega_{d})^{-1/(d+1)}|A|e^{-\beta/(d+1)}).$ (2.11)
Also as $L\to\infty$,
$\mathbb{P}\left[\omega_{d}\tilde{\tau}_{L}^{d+1}-d(d+1)\log L-d^{2}\log\log L\leq\beta\right]$
$\to\exp\Big{(}-c_{d}\Big{(}\frac{(d+1)^{d^{2}}}{d^{d}\omega_{d}}\Big{)}^{1/(d+1)}|A|e^{-\beta/(d+1)}\Big{)}.$ (2.12)
In Appendix A, we shall verify that this result is consistent with the case $\gamma=v=1$ of [1, Theorem 4]. The proof in [1] refers to an unpublished preprint; for completeness we shall present the (fairly short) proof of Proposition 2.8 in Section 4.
Our main new results in this section are concerned with $T_{\rho}$ and $\tau_{L}$, the cover times for the restricted J-M model. We first give a result for the case where $A$ is a polygon in $d=2$.
@@
Suppose that $d=2$ and $A$ is polygonal. Let $\beta\in\mathbb{R}$. Then as $\rho\to\infty$,
$\mathbb{P}[\pi\rho T_{\rho}^{3}-2\log\rho-4\log\log\rho\leq\beta]\to\exp(-(4\pi)^{-1/3}|A|e^{-\beta/3}-(2\pi^{2})^{-1/3}|\partial A|e^{-\beta/6}).$ (2.13)
Also as $L\to\infty$,
$\mathbb{P}[\pi\tau_{L}^{3}-6\log L-4\log\log L\leq\beta]\to\exp\Big{(}-\Big{(}\frac{81}{4\pi}\Big{)}^{1/3}|A|e^{-\beta/3}-\Big{(}\frac{9}{2\pi^{2}}\Big{)}^{1/3}|\partial A|e^{-\beta/6}\Big{)}.$ (2.14)
Our result for general $d$ and $A$ with $C^{2}$ boundary involves the constant $c^{\prime}_{d}$ defined by
$c^{\prime}_{d}$ $:=\frac{c_{d-1}\omega_{d-1}^{2d-3}}{\omega_{d-2}^{d-1}d^{d-1}}\Big{(}\frac{(d-1)^{d}}{2\omega_{d}^{d}}\Big{)}^{(d-1)/(d+1)}.$ (2.15)
@@
Suppose that $d\geq 2$ and $\partial A\in C^{2}$ with $\overline{A^{o}}=A$. Let $\beta\in\mathbb{R}$. Then as $\rho\to\infty$, if $d=2$ then (2.13) holds while if $d\geq 3$ then
$\mathbb{P}[\omega_{d}\rho T_{\rho}^{d+1}-2(d-1)\log\rho-2d(d-1)\log\log\rho\leq\beta]\to\exp(-c^{\prime}_{d}|\partial A|e^{-\beta/(2d+2)}).$ (2.16)
Also as $L\to\infty$, if $d=2$ then (2.14) holds while if $d\geq 3$ then
$\mathbb{P}[\omega_{d}\tau_{L}^{d+1}-2(d^{2}-1)\log L-2d(d-1)\log\log L\leq\beta]$
$\to\exp(-(d+1)^{d(d-1)/(d+1)}c^{\prime}_{d}|\partial A|e^{-\beta/(2d+2)}).$ (2.17)
Figure 1: Empirical distributions of the standardised coverage times $T_{\rho}$ and $\tilde{T}_{\rho}$ as described in Proposition 2.8 and Theorem 2.9, obtained by sampling many independent realisations of the Johnson-Mehl process in $[0,1]^{2}$. The standardisation function $g(T,\rho)=\pi\rho T^{3}-2\log\rho-4\log\log\rho$ is the same for both $T_{\rho}$ and $\tilde{T}_{\rho}$. In [1], Chiu proves a weak limit theorem for $\tilde{T}_{\rho}$ (or rather, for $\tilde{\tau}_{L}$, which is equivalent to a limit theorem for $\tilde{T}_{\rho}$) with a different standardisation. In the large-$\rho$ limit, we show in Appendix A that our weak law in Proposition 2.8 is consistent with [1]. Chiu suggested that the same limit theorem should apply to $\tau_{L}$ as to $\tilde{\tau}_{L}$, as the boundary effects would not affect the limit. Our Theorem 2.9, and the above figure, show that in fact the boundary effects do affect the limiting distribution of $g(T_{\rho},\rho)$.
In short, when $d=2$, for $\rho$ large the last point of $A$ to get covered in the restricted J-M model could lie either on the boundary or in the interior of $A$, while for $d\geq 3$, when $\rho$ is large the last point to get covered is very likely to lie on the boundary of $A$.

### v3

@@ Abstract
Given a compact planar region $A$, let $\tau_{A}$ be the (random) time it takes for the Johnson-Mehl tessellation of $A$ to be complete, i.e. the time for $A$ to be fully covered by a spatial birth-growth process in $A$ with seeds arriving as a unit-intensity Poisson point process in $A\times[0,\infty)$, where upon arrival each seed grows at unit rate in all directions. We show that if $\partial A$ is smooth or polygonal then $\mathbb{P}[\pi\tau_{sA}^{3}-6\log s-4\log\log s\leq x]$ tends to $\exp(-(\frac{81}{4\pi})^{1/3}|A|e^{-x/3}\linebreak-(\frac{9}{2\pi^{2}})^{1/3}|\partial A|e^{-x/6})$ in the large-$s$ limit; the second term in the exponent is due to boundary effects, the importance of which was not recognized in earlier work on this model. We present similar results in higher dimensions (where boundary effects dominate). These results are derived using new results on the asymptotic probability of covering $A$ with a high-intensity spherical Poisson Boolean model restricted to $A$ with grains having iid small random radii, which generalize recent work of the first author that dealt only with grains of deterministic radius.
@@ 1 Introduction
This model dates back at least to work of Kolmogorov and of Johnson and Mehl in the 1930s on modelling crystallization processes; other applications include growth of surface film on metals, and more recent applications in neurobiology are described in [4]. For further discussion and references, see also [10, 11].
A natural variant is the restricted Johnson-Mehl tessellation, which we define as follows. Given a specified compact region $A\subset\mathbb{R}^{d}$, we allow only for seeds that arrive inside the region $A$, that is, the seeds are generated by a Poisson process in $A\times[0,\infty)$. Cells grow by the same rules as described before, and the restrictions of the cells to $A$ ultimately tessellate $A$, as shown in Figure 1.
@@
In [3, page 893], Chiu rightly distinguishes between the J-M and restricted J-M models. However, he goes on to assert that ‘all theorems in this paper are valid for both models’ because ‘almost surely edge effects do not play a role in the limiting behaviour of $T_{L}$’ (Chiu’s $T_{L}$ is equivalent to our $\tau_{L}$). In other words, Chiu seems to be asserting that $\tau_{L}$ and $\tilde{\tau}_{L}$ have the same limiting distribution.
It is our contention that this assertion is incorrect. In dimensions $d\geq 2$, the limiting behaviour of $T_{\rho}$ is different from that of $\tilde{T}_{\rho}$, and the limiting behaviour of $\tau_{L}$ is different from that of $\tilde{\tau}_{L}$; in other words, edge effects do play a role. We justify this assertion with results identifying the limiting distribution of $T_{\rho}$ (suitably scaled and centred) and showing it is different from that of $\tilde{T}_{\rho}$. Likewise we show that the limiting distribution of $\tau_{L}$ (suitably scaled and centred) is different from that of $\tilde{\tau}_{L}$; our limiting result for $\tilde{\tau}_{L}$ is consistent with the result in [3]. Our results for $T_{\rho}$ and $\tau_{L}$ apply when $A$ is polygonal in $d=2$, or when it has a smooth (more precisely, $C^{2}$) boundary for general $d\geq 2$.
When $A=[0,1]^{d}$, we do not provide detailed limiting distributions for $T_{\rho}$ and $\tau_{L}$ except in the case $d=2$. To give a detailed limit distribution when $d\geq 3$ would require careful consideration of all faces of all dimensions and is beyond the scope of this paper. However, the time to cover all the (d1)(d-1)-dimensional faces will be a lower bound for the actual cover time so boundary effects can certainly not be neglected in this case either.
@@
The Johnson-Mehl cover time is the maximum of the random field $(\Xi_{x},x\in A)$, where $\Xi_{x}$ denotes the the time at which at which $x$ is covered. For the restricted J-M model, this maximum will be achieved at a vertex $x$ of the restricted Johnson-Mehl tessellation of $A$, as illustrated in Figure 1. (In fact there are other possibilities; e.g. if $A$ is polygonal with a sharp corner then $\sup_{x\in A}\Xi_{x}$ could be achieved at a corner of $A$ but these other possibilities have vanishing probability in our limit regimes, at least when $d=2$ or $d=3$.) Thus it is the maximum of a large finite random number of (perhaps weakly) dependent variables, and one might perhaps expect one of the classical extreme value distributions such as the Gumbel to arise in the limiting regimes that we consider.
It turns out that the limit distribution for the restricted model is indeed Gumbel in 3 or more dimensions, but in two dimensions it is a two-component extreme value distribution (see Remark 2.8 below). In short, this is because for $d=2$, the maximum could be achieved either at a vertex in the interior of $A$, or at a vertex on the boundary of $A$, and these different possible contributions scale differently as $\rho\to\infty$; in higher dimensions, the maximum is very likely to be at the boundary.
For the restricted SPBM, again the probability of coverage can be framed as the extreme value of a random field. Label the Poisson points in $A$ as $\{p_{1},\ldots,p_{N}\}$ and let $Y_{i}$ be the random radius associated with Poisson point $p_{i}$. Define the coverage threshold $R$ to be the smallest $r$ such that the union of balls of radius $rY_{i}$ centred on $p_{i}$ covers $A$. Then for any scaling factor $r>0$, the probability that that $A$ is fully covered by balls of radius $rY_{i}$ equals the cdf $F_{R}(r):=\mathbb{P}[R\leq r]$. The threshold is the maximum of a random field, again denoted $(\Xi_{x},x\in A)$, where now $\Xi_{x}$ is the smallest $r$ such that $x\in\cup_{i}B(p_{i},rY_{i})$. Moreover, once again for high intensities the maximum will typically (at least in low dimensions) be achieved at a vertex of a certain tessellation of $A$, namely the division of $A$ into cells $C_{i},1\leq i\leq N$ with
$C_{i}:=\big\{y\in A:\frac{\|y-p_{i}\|}{Y_{i}}\leq\frac{\|y-p_{j}\|}{Y_{j}}~\forall~j\in\{1,\ldots,N\}\big\},$
where $\|\cdot\|$ denotes the Euclidean norm on $\mathbb{R}^{d}$; here the cells might not be connected. See Figure 1 for an illustration (which does indeed include a disconnected cell) and Remark 2.15 for further discussion (again the cells have a dynamical interpretation). In the special case where the $Y_{i}$ are a deterministic constant this is simply the Voronoi tessellation of $A$ induced by the Poisson process in $A$. In fact, in this case the coverage threshold is the largest circumscribed radius of the Poisson-Voronoi tessellation of $A$, as discussed in [15] for the restricted SPBM and in [2] for the unrestricted SPBM.
Figure 1: Tessellations of a disc AA (diameter $0.9$) by the restricted Johnson–Mehl process with $\rho\approx 125$ (left) and restricted spherical Poisson Boolean model with $Y_{i}$ exponentially distributed (right). In both cases the seeds are marked inside each cell with a red dot, and $\mathrm{argmax}_{x\in A}\Xi_{x}$ (the “last location covered”) is marked with a blue square. Later-arriving seeds are marked with paler dots (left diagram); larger dots indicate larger $Y_{i}$ (right diagram). All vertices of both tessellations are of degree 3, including those on the boundary.
Thus in all cases we are interested in the distribution of a random variable (either the cover time or the coverage threshold) given by the maximum of a certain geometrically-defined random field on $A$. Other somewhat related topics within this genre of geometric extreme value theory (as it might reasonably be termed) include multivariate scan statistics and clique number of random geometric graphs (see e.g. [1], [14]), and largest nearest $k$-neighbour link ($k$-NNL) of a random sample of points (see [16] and references therein in a Euclidean space, [12] in hyperbolic space).
@@ 2 Statement of results
For $x\in\mathbb{R}^{d}$ and $r>0$ set $B(x,r):=\{y\in\mathbb{R}^{d}:\|y-x\|\leq r\}$. (We write $B_{(d)}(x,r)$ for this if we wish to emphasise the dimension.) For $D\subset\mathbb{R}^{d}$, let $\overline{D}$ and $D^{o}$ denote the closure of $D$ and interior of $D$, respectively, and set $\partial D:=\overline{D}\setminus D^{o}$, the topological boundary of $D$. For $r>0$, let $D^{(r)}:=\{x\in D:B(x,r)\subset D^{o}\}$, the ‘$r$-interior’ of $D$. Let $|D|$ denote the Lebesgue measure (volume) of $D$, and $|\partial D|$ the perimeter of $D$, i.e. the (d1)(d-1)-dimensional Hausdorff measure of $\partial D$, when these are defined. Write $\log\log t$ for $\log(\log t)$, $t>1$.
@@
We say that $D$ has $C^{2}$ boundary (for short: $\partial D\in C^{2}$) if for each $x\in\partial D$ there exists a neighbourhood $U$ of $x$ and a real-valued function $f$ that is defined on an open set in $\mathbb{R}^{d-1}$ and twice continuously differentiable, such that $\partial D\cap U$, after a rotation, is the graph of the function $f$. We say that $\partial D\in C^{1,1}$ (a weaker condition) if for each $x$ the function $f$ satisfies only that $f$ is continuously differentiable with Lipschitz first order partial derivatives.
@@
$c_{d}:=\frac{1}{d!}\left(\frac{\sqrt{\pi}\;\Gamma(1+d/2)}{\Gamma((d+1)/2)}\right)^{d-1}.$ (1)
Note that $c_{1}=c_{2}=1$, and $c_{3}=3\pi^{2}/32$. Moreover, using Stirling’s formula one can show that $c_{d}^{1/d}\sim e(\pi/(2d))^{1/2}$ as $d\to\infty$. The constant $c_{d}$ is denoted $\psi_{d}$ in [3, page 894], while $\omega_{d}$ is denoted $\theta_{d}$ in [15]. Later it will be useful to have $\omega_{0}$ defined: we set $\omega_{0}:=1$.
@@ 2.1 Results for the Johnson-Mehl model
$T_{\rho}:=\inf\left\{t>0:A\subset\cup_{(x,s)\in{\cal H}_{\rho}\cap(A\times[0,t])}B(x,t-s)\right\};$ (2)
$\tilde{T}_{\rho}:=\inf\left\{t>0:A\subset\cup_{(x,s)\in{\cal H}_{\rho}\cap(\mathbb{R}^{d}\times[0,t])}B(x,t-s)\right\}.$ (3)
@@
$\tau_{L}:=\inf\left\{t>0:A_{L}\subset\cup_{(x,s)\in{\cal H}_{1}\cap(A_{L}\times[0,t])}B(x,t-s)\right\};$ (4)
$\tilde{\tau}_{L}:=\inf\left\{t>0:A_{L}\subset\cup_{(x,s)\in{\cal H}_{1}\cap(\mathbb{R}^{d}\times[0,t])}B(x,t-s)\right\}.$ (5)
@@
We are concerned with the asymptotic distribution of the cover times $T_{\rho}$ and $\tilde{T}_{\rho}$ as $\rho\to\infty$, and the asymptotic distribution of $\tau_{L}$ and $\tilde{\tau}_{L}$ as $L\to\infty$. We start with a result in general $d$ for the unrestricted cover time $\tilde{T}_{\rho}$ (and for $\tilde{\tau}_{L})$. This is simpler to deal with than $T_{\rho}$ because boundary effects are avoided.
@@
Suppose $A\subset\mathbb{R}^{d}$ is compact and Riemann measurable with $|A|>0$. Let $\beta\in\mathbb{R}$. Then
$\lim_{\rho\to\infty}\mathbb{P}\left[\omega_{d}\rho\tilde{T}_{\rho}^{d+1}-d\log\rho-d^{2}\log\log\rho\leq\beta\right]$
$=\exp(-c_{d}(d^{d}\omega_{d})^{-1/(d+1)}|A|e^{-\beta/(d+1)})$ (6)
$=\lim_{L\to\infty}\mathbb{P}\left[\omega_{d}\tilde{\tau}_{L}^{d+1}-d(d+1)\log L-d^{2}\log\log L-d^{2}\log(d+1)\leq\beta\right].$ (7)
In Appendix A, we shall verify that this result is consistent with the case $\gamma=v=1$ of [3, Theorem 4]. The proof in [3] refers to an unpublished preprint; for completeness we shall present the (fairly short) proof of Proposition 2.1 in Section 3.
Our main new results in this subsection are concerned with $T_{\rho}$ and $\tau_{L}$, the cover times for the restricted J-M model. We first give a result for the case where $A$ is a polygon in $d=2$.
@@
Suppose that $d=2$ and $A$ is polygonal. Let $\beta\in\mathbb{R}$. Then
$\lim_{\rho\to\infty}\mathbb{P}[\pi\rho T_{\rho}^{3}-2\log\rho-4\log\log\rho\leq\beta]$
$=\exp(-(4\pi)^{-1/3}|A|e^{-\beta/3}-(2\pi^{2})^{-1/3}|\partial A|e^{-\beta/6})$ (8)
$=\lim_{L\to\infty}\mathbb{P}[\pi\tau_{L}^{3}-6\log L-4\log\log L-\log 81\leq\beta].$ (9)
Our result for general $d$ and $A$ with $C^{2}$ boundary involves a constant $c^{\prime}_{d}$ defined by
$c^{\prime}_{d}$ $:=\frac{c_{d-1}\omega_{d-1}^{2d-3}}{\omega_{d-2}^{d-1}d^{d-1}}\Big(\frac{(d-1)^{d}}{2\omega_{d}^{d}}\Big)^{(d-1)/(d+1)}.$ (10)
@@
Suppose that $d\geq 2$ and $\partial A\in C^{2}$ with $\overline{A^{o}}=A$. Let $\beta\in\mathbb{R}$. If $d=2$ then (8) and (9) hold while if $d\geq 3$ then
$\lim_{\rho\to\infty}\mathbb{P}[\omega_{d}\rho T_{\rho}^{d+1}-2(d-1)\log\rho-2d(d-1)\log\log\rho\leq\beta]$
$=\exp(-c^{\prime}_{d}|\partial A|e^{-\beta/(2d+2)})$ (11)
$=\lim_{L\to\infty}\mathbb{P}[\omega_{d}\tau_{L}^{d+1}-2(d^{2}-1)\log L-2d(d-1)\log\log L$
$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~-2d(d-1)\log(d+1)\leq\beta].$ (12)
Figure 2: Empirical distributions of the standardised coverage times $T_{\rho}$ and $\tilde{T}_{\rho}$ as described in Proposition 2.1 and Theorem 2.2, obtained by sampling many independent realisations of the Johnson-Mehl process in $0,1]^{2}$. The standardisation function $g(T,\rho)=\pi\rho T^{3}-2\log\rho-4\log\log\rho$ is the same for both $T_{\rho}$ and $\tilde{T}_{\rho}$.
Let $\mathsf{Gu}$ denote a standard Gumbel random variable, i.e. one with cdf $\mathbb{P}[{\mathsf{Gu}}\leq\beta]=\exp(-e^{-\beta}),\beta\in\mathbb{R}$. The limit in (6) is the cdf of the random variable
$(d+1){\mathsf{Gu}}+\log(c_{d}^{d+1}d^{-d}/\omega_{d})+(d+1)\log|A|,$
so Proposition 2.1 says that $\omega_{d}\rho\tilde{T}_{\rho}^{d+1}$ suitably centred by a constant depending on $\rho$ and $d$ but not $A$, and $\omega_{d}\tilde{\tau}_{L}^{d+1}$ suitably centred likewise, both converge in distribution to the random variable $(d+1)({\mathsf{Gu}}+\log|A|)$.
Similarly, Theorem 2.3 says for $d\geq 3$ that $\omega_{d}\rho T_{\rho}^{d+1}$ suitably centred by a constant depending on $\rho$ and $d$ but not $A$, converges in distribution to $(2d+2)({\mathsf{Gu}}+\log|\partial A|)$, as does $\omega_{d}\tau_{L}^{d+1}$.
Theorem 2.2 says for $d=2$ that $\pi\rho T_{\rho}^{3}-2\log\rho-4\log\log\rho$, and $\pi\tau_{L}^{3}-6\log L-4\log\log L-\log 81$, both converge in distribution to the random variable
$\max(3({\mathsf{Gu}}+\log|A|)-\log(4\pi),6({\mathsf{Gu}}^{\prime}+\log|\partial A|)-\log 4\pi^{4}),$ (13)
where $\mathsf{Gu}$$^{\prime}$ is another standard Gumbel variable, independent of $\mathsf{Gu}$}.